OpenStudy (dengeki_daisy):

Which would have a higher boiling point? a.1.0 M KNO3 b.1.5 M KNO3 c.0.75M KNO3 d.2.0M KNO3

OpenStudy (dengeki_daisy):

does anybody know

OpenStudy (cuanchi):

what is your guess and why?

OpenStudy (dengeki_daisy):

Well I don't know how to work it out....

OpenStudy (cuanchi):

do you know what is boiling point elevation?

OpenStudy (dengeki_daisy):

I could tell you but I don' know what it means

OpenStudy (cuanchi):

tell me

OpenStudy (cuanchi):

d.2.0M KNO3

OpenStudy (dengeki_daisy):

yeah I figured that was it

OpenStudy (dengeki_daisy):

delta T = i*Kb*M i - 2 ]Kb = 1.52 M = 1 M Then delta T = 2*1.52*1 = ? dT = 3.04 for a.

OpenStudy (cuanchi):

great! good luck you are a genius!!

OpenStudy (dengeki_daisy):

thank you!